OpenStudy (anonymous):

can you help me to differentiate (x+y)^3 = x^2 + y ?

OpenStudy (turingtest):

implicitly yeah?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

3(x+y)^2(1+y')=2x+y' 3(x+y)^2-2x=y'-3(x+y)^2y'=(1-3(x+y)^2)y' \[y'={3(x+y)^2-2x \over1-3(x+y)^2}\]expand and simplify